### v1

# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

Abstract.
Performance plays a critical role in ensuring the smooth operation of any mobile application, directly influencing user engagement and retention. Android applications are no exception. However, unlike functionality issues, performance issues are more challenging to discover as their root causes are sophisticated and typically emerge under specific payloads. To tackle this problem, researchers have dedicated substantial efforts to proposing automatic approaches for understanding, detecting, and resolving performance issues. Despite these endeavors, it still remains unknown what the status quo of Android performance analysis is, and whether existing approaches can indeed accurately reflect real performance issues. To fill this research gap, we conducted a systematic literature review followed by an explanatory study to explore relevant studies and real-world challenges. Our findings reveal that current tools have limited capabilities, covering only 17.50% of the performance issues. Additionally, existing datasets encompass only 27.50% of the issues and are very limited in size. We also show real-world issue patterns, underscoring the huge gap between the identified techniques and practical concerns. Furthermore, possible solutions are provided to guide future research towards achieving effective performance issue detection and resolution.
conference: ; ; copyright: none
1. Introduction
With the increasing popularity of smartphones, the Android system has become one of the largest mobile platforms in the world, holding a 70.69% market share worldwide (iPhoneAndroidStatistics, ). The exponential growth of Android usage has seen an unprecedented expansion within the mobile app industry, with millions competing for users’ attention. In such a fiercely competitive circumstance, developers are under pressure to create new, fancy apps in the quickest manner. However, this emphasis on rapid development can sometimes come at the expense of app quality, particularly in terms of performance.
Unlike functionality issues that directly cause app malfunctions or failures, performance issues occur when the app functions correctly but its efficiency is low, resource usage is unreasonable, or user interaction is poor, significantly reducing user satisfaction and software reliability (Li2019DetectingAD, ; Liu2014CharacterizingAD, ; Gao2017EveryPC, ; Liu2018NavyDroidAE, ). These issues include, but are not limited to, excessive energy consumption (Banerjee2018EnergyPatchRR, ; Pathak2012WhatIK, ; Carette2017InvestigatingTE, ), resource leaks (Wu2016LightWeightIA, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Yan2013SystematicTF, ), and responsiveness delays (Farooq2018RuntimeDroidRR, ; Kang2016DiagDroidAP, ; Lin2024AgingOG, ). For example, Liu et al. (Liu2014CharacterizingAD, ) demonstrate that wasteful energy consumption can occur in background threads that continue to update invisible GUIs or perform complex network computations, rapidly depleting battery life and degrading app performance. Similarly, Yan et al. (Yan2013SystematicTF, ) reveal resource leak defects associated with events such as screen rotation, app switching, pressing the back button, and file or database access.
However, due to the non-functional nature of performance issues, their problems do not immediately manifest. Additionally, performance issues can be caused by multiple factors, requiring in-depth analysis and debugging to diagnose and identify the root causes (Liu2016FixingRL, ; Habchi2018OnAL, ; Jabbarvand2019SearchBasedET, ; Li2019CharacterizingAD, ; Liu2014GreenDroidAD, ). Furthermore, some performance issues only appear under specific usage scenarios or loads (Behrouz2016EnergyawareTM, ; Jabbarvand2019SearchBasedET, ; Yan2013SystematicTF, ; Farooq2018RuntimeDroidRR, ). Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the testing phase. Therefore, it is crucial to use effective automated tools to analyze, detect, and fix performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ).
Unfortunately, at the current stage, researchers face difficulties in designing automated tools to comprehensively address Android performance issues due to: 1) Lack of clear classification of performance issues and their root causes, making it hard to fully understand the performance issues; 2) Uncertainty about the state quo of existing approaches, making it unclear to what extent these tools are reliable for detecting or fixing performance issues; 3) The absence of benchmark datasets in evaluating state-of-the-art tools; 4) Uncertainty about whether state-of-the-art tools and datasets accurately reflect real performance issues.
To guide future research, we first conducted a systematic literature review to provide a holistic taxonomy for existing performance issues and their underlying root causes. We then identified research gaps between state-of-the-art tools and datasets relative to the targeted research objectives (i.e., performance issues). Subsequently, we analyzed real-world performance issues to pinpoint differences between existing literature and practical challenges. Through these investigations, our goal is to offer a thorough understanding of Android performance issues and to guide the development of more effective automated tools and comprehensive datasets for evaluation. In this paper, we aim to explore the following research questions:
RQ1: What is the status quo of Android performance issues analysis in research field?
  • What performance issues are being targeted by researchers? The lack of a universally accepted taxonomy leads to many papers targeting the same performance problems using different terms. Additionally, since one factor can lead to multiple performance issues and one issue can be caused by multiple factors, creating a clear classification system is challenging. Our work filled this gap by distinguishing between the consequences and factors of performance issues. We identified five categories of Android performance issues and 21 distinct root causes.
  • What approaches have been developed for performance issues? To solve the above performance issues, researchers have proposed various kinds of techniques. Thus, in the second research question, we aim to understand the status quo of current techniques, seeking to comprehensively grasp the research gap between performance issues and automatic tools. In total, we classified them into three primary techniques: static-based, dynamic-based, and deep learning-based approaches. Static analysis is the most frequently adopted technique, while deep learning-based methods are much less commonly utilized for performance issues. Moreover, it is surprising to find that only 7 out of 21 factors are targeted by these tools. This evidence confirms a significant gap between performance issues and automated resolutions. Additionally, the majority of these tools focus on issue detection rather than fixing, highlighting another research gap concerning the need for proactive problem-solving.
  • What datasets have been used in Android performance analysis? High-quality datasets are essential for evaluating performance analysis tools. We found that current datasets cover only 11 out of 21 factors, addressing a limited range of issues. Furthermore, the number of existing datasets is very small. This highlights an urgent need for a comprehensive public dataset.
RQ2: What are gaps between cutting-edge studies and real-world scenarios?
  • What real-world performance issues in the wild? Given that performance is an inherently practical concern, we proceeded to conduct a deeper analysis of real-world issues to pinpoint differences between existing literature and practical challenges. Our examination revealed one previously unrecognized performance issue and 19 previously unacknowledged factors across real-world platforms. This suggests that numerous factors contributing to performance issues in practical settings have not received attention in research. Furthermore, we uncovered a misalignment between the primary concerns of researchers, developers, and users. These discoveries underscore a substantial gap in fully comprehending and mitigating performance issues in real-world scenarios.
  • What are the common patterns exhibited in real-world performance issues? Based on our systematic analysis for real-world data, we identified six common code patterns: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns. These newly discovered patterns provide insights for designing automated tools to address performance issues and for creating datasets to evaluate the effectiveness of these tools.
2. Studies on Performance Issues
In this work, we are interested in how far we are towards automating performance issues analysis in Android apps. Thus, in this section, we first resort to a systematic literature review (SLR) to understand the status quo of Android app performance analysis.
2.1. Literature Review
To conduct systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 1 illustrates the working process which comprises 5 steps.
2.1.1. Keywords Identification
To investigate the status quo of Android performance analysis, we identified a bunch of search keywords to locate relevant publications on this topic. Specifically, the keywords comprise two groups (i.e., G1 AND G2). Each group is made up of several keywords. The search string is then constructed by combining them, i.e., g1 AND g2, where g1 and g2 represent the disjunction of keywords from groups G1 and G2, respectively. Please note that these keywords are based on our initial understanding of performance issues and are intended to help us identify “seed” papers. As we review these papers, we could uncover new performance issues, prompting a snowballing process that could lead us to further relevant studies. The * symbol represents any string of characters. For example, *phone* can refer to smartphone, phones, or phone. Similarly, issue* can represent issues or issue. Keywords are as follows:
G1: android, mobile, *phone*
G2: performance, resource, energy, responsiveness, issue*
Figure 1. The working process of our systematic literature review.
Table 1. CORE A/A* ranked software engineering and mobile venues.
Type Source Research Codes Venues
Journals CORE2020 0803, 0805 TSE, TOSEM, JSS, IST, TMC
Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, ISCA, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
2.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library.
From this initial search, we retrieved a total of 10,766 papers. However, not all of them were of high quality, and many were outside our area of focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, to further guarantee the quality of papers, we exclusively chose those published from conferences rated CORE A*/A  [^1] CORE provides conference and journal ratings. https://www.core.edu.au/home. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the title of relevant publications. As a result, we successfully identified 274 publications distributed across 22 venues as listed in Table 1 (i.e., there is no relevant paper identified in 4 of the venues).
2.1.3. Paper Exclusion
To gather relevant papers as possible, we initially considered all returned results from our search. However, not every paper pertains to Android performance issues. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
  1. (1)
    Short papers (i.e., less than six pages in double-column format or 11 pages in single-column format) were excluded.
  2. (2)
    Papers targeting on non-Android mobile devices (e.g., Windows Phone, IOS devices) were excluded.
  3. (3)
    Papers using Android performance issues as metrics/indicators for comparison were excluded. For example, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations.
  4. (4)
    Papers targeting on Android performance issues but not focused on the mobile app level were excluded. For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope focused on mobile app issues stemming from poor coding practices by developers.
After applying these exclusion criteria, we retained 27 papers that are closely related to the topic of performance issues analysis at Android mobile level.
2.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 27 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 48 papers. Please note that during snowballing, we also collected 82 keywords related to performance issues for real-world data collections in Section 3.1.2.
2.1.5. Cross Check
Even with the relevant selection criteria, it is not guaranteed that the quality of all identified papers is desirable. Therefore, we evaluated the quality of each paper by reading its full text. Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 48 papers to answer our research questions defined in Section 1.
2.2. Result
In the following paragraphs of Section 2, we summarize the state quo from thorough reading of these papers and attempt to understand: 1) Targeted performance issues, which involve key factors that lead to problems needing resolution, 2) Automatic techniques designed to address mobile performance issues, and 3) Datasets designed to evaluate the effectiveness of these automatic techniques.
Table 2. The top problems targeted by the examined papers.
Category Factor Problem Example
Energy Consumption Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, )
Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, )
Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, )
Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
UI Rendering Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, )
View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Responsiveness Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image 1) Decoding large images in the UI thread, 2) Loading large image, 3) dynamically adjusting content during activity-switching events (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Dynamic UI Loading Frequent layout inflations during scrolling  (Liu2014CharacterizingAD, )
UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
SQL Statement Retrieving more columns/rows than necessary in SQL queries  (Lyu2019QuantifyingTP, )
Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, )
Activity Loading new resources during activity restarts may freeze the UI  (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Processing tasks in a single-threaded pool of AsyncTask since Android 3.0 (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, )
Memory Consumption Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image Decoding images without resizing; leaking image objects  (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Service Failing to destroy a dead service  (Song2019ServDroidDS, )
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
AsyncTask Failing to destroy AsyncTask when it holds active GUI components (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Fan2018EfficientlyMA, )
View Making allocations inside onDraw() routines  (Habchi2019TheRO, ; Couto2020EnergyRF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Data Loss Activity Failing to save and restore GUI data during activity restarts from runtime changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Updating destroyed GUI components when it has been recreated using AsyncTask (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Distorted UI Display Activity Overlapping fragments occur when a new activity is restarted with a new fragment attached while the old fragment is still retained by the system (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
2.2.1. Result - Targeted Problem
We now highlight the representative problems targeted by our fellow researchers (cf. Table 2). Based on the literature review results, we manually summarized and identified five major categories of Android mobile performance issues: Energy Consumption, Memory Consumption, Responsiveness, Data Loss, and Distorted UI Display. Then we reviewed the root causes (a.k.a factors) for each category thoroughly and demonstrate them in Table 2 along with the specific performance problems our researcher fellows are working on.
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The lack of a universally accepted taxonomy results in many papers targeting the same performance problems but using different descriptive terms. Additionally, since one factor can lead to multiple performance issues (consequences), and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. Therefore, our work addresses this gap by distinguishing between the consequences and factors of performance issues. Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors. To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner. We now elaborate on the five major categories, each with the corresponding factors along with an example presented in Table 2.
Energy Consumption. As shown in Fig. 2, energy consumption receives the most research attention, with 34 out of the examined papers (70.83%) and its increasing trend is expected to continue. This result is not surprising given that the most urgent task at the moment is to ensure that the app remains user-friendly and sustainable in the rapidly evolving app market, which is why many previous studies focused on this research topic. Specifically, we found that most of the energy consumption issues stem from a common pattern: frequent energy-consuming operations occurring together with loops/lifecycle functions. For example, Lyu et al. (Lyu2018RemoveRF, ) found that database writes within loops trigger inefficient autocommit behaviors, leading to significant energy waste. Liu et al. (Liu2014CharacterizingAD, ) discovered that even when the GUI is in the background and not visible, threads continue updating the UI, causing unnecessary energy waste. In addition, invoking APIs for requiring resources (e.g., wakelock, sensor) when they are not needed can also results in a substantial increase in energy consumption (Liu2014GreenDroidAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Banerjee2018EnergyPatchRR, ).
Responsiveness. Another important aspect of Android performance issues is responsiveness. Fig. 2 shows that 18 reviewed articles (37.50%) are crafted for poor responsiveness, leading to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. Specifically, we identified 10 distinct factors that contribute to this category. Most primary studies (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ) discovered that heavy operations associated with UI threads can slow responsiveness. Additionally, Gao et al. (Gao2017EveryPC, ) and Sarra et al. (Habchi2019TheRO, ) show that improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
Memory Consumption. There are 14 papers (29.17%) related to memory consumption issues in the collected studies. These memory consumption issues have diverse factors, often stemming from inefficient coding that leads to excessive memory use. For example, decoding images without resizing them results in extensive memory consumption (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ). Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks (Lin2015StudyAR, ; Pan2020StaticAC, ).
Data Loss With regard to the data loss problem, Fig. 2 indicates that 7 papers (14.58%) attempt to analyze data loss issues, which significantly degrade user experience when data fails to be saved or restored in UI threads. For instance, studies have revealed that failing to save and restore GUI data during activity restarts can cause existing data to vanish instantly when the screen orientation changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ).
Distorted UI Display. Another aspect of performance issues relates to distorted UI display, which consists of only 2.08% (i.e., 1 study) of the collected studies. Specifically, Umar et al. (Farooq2018RuntimeDroidRR, ) shows that overlapping fragments occur when initializing a new fragment while retaining the old one, causing a poor user experience.
Figure 2. Annual publications on each performance issue.
Discussion. Among all the collected studies on performance issues for Android, it appears that energy consumption is the most frequently studied topic, followed by responsiveness analysis. Additionally, it is noteworthy that an increasing number of recent studies have started to address specific issues related to memory consumption and data loss (e.g., runtime change handling (Farooq2018RuntimeDroidRR, ; Chen2023TransparentRC, )), although the total number of relevant studies remains small. Moreover, the amount of literature focusing on distorted UI display is relatively small, indicating a need for further in-depth research. While this review categorizes the existing literature’s research objectives into five categories, the scope of Android performance issues can be much broader. Therefore, future research should go beyond these categories and propose new approaches tailored to emerging app scenarios.
Answers to RQ1.1: Current research studies focus on five categories of performance issues, prioritized by their prevalence: Energy Consumption (70.83%), Responsiveness (37.50%), Memory Consumption (29.17%), Data Loss (14.58%), and Distorted UI Display (2.08%). We provide a holistic taxonomy of Android performance issues, encompassing 5 major categories and 21 distinct root causes. Performance issues are sophisiticated, as a single issue may stem from multiple factors and a single factor can cause multiple issues.
2.2.2. Result - Technique
To address the above performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the the list of available automatic approaches collected from our examined papers. Despite numerous claims of proposed automated tools, only a small fraction (i.e., 10 out of the 48 examined papers (20.83%)) works have made their tool publicly available. This indicates a significant gap between automatic tools with their practical usage to within our community. In addition, among the identified factors leading to performance issues, only 7 out of 21 are targeted by existing tools. The fact that the majority of the factors are overlooked by SOTA tools indicates a huge gap in addressing performance issues. Moreover, most automatic tools are designed with a primary focus on simply detecting performance issues rather than actively fixing them. This emphasizes a need for a shift towards developing tools that not only identify but also offer solutions to enhance performance. To better understand how these available tools are implemented, we now summarize the techniques adopted in these approaches.
Table 3. A summary of automatic approaches used for Android mobile performance analysis. Static denotes Static Analysis, Dynamic denotes Dynamic Analysis, and DL denotes Deep Learning methods.
Tool/Reference Approach Technique Stage Target Factor Link
DDLDroid (Zhou2023DDLDroidED, ) Static Reverse Engineering Taint analysis Detect Activity  (DDLDroidTool, )
DLD (Riganelli2020DataLD, ) Dynamic Runtime Monitoring Model-based Testing Detect Activity  (Riganelli2020DataLDTool, )
iFixDataloss (Guo2022DetectingAF, ) Static Dynamic Fuzzing Search-based Testing Static Program Repair Detect Fix Activity  (Guo2022DetectingAFTool, )
RCHDroid (Chen2023TransparentRC, ) Dynamic Runtime Management Garbage Collection Fix Activity  (Chen2023TransparentRCTool, )
VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Activity Resource  (Lu2022DetectingRUTool, )
ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Database Loop  (Lyu2018RemoveRFTool, )
ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural- Code Analysis Detect Service  (Song2019ServDroidDSTool, )
Static Analysis. Static approaches are the art of reasoning about the behavior of programs without actually running them. It has been widely used (70.00%) to either detect or fix performance issues, as shown in Table 3. Specifically, taint analysis, symbolic execution, code pattern analysis and data flow analysis are the most commonly adopted techniques for rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Guo2022DetectingAF, ).
Dynamic Analysis. There are only four (i.e. 40.00%) primary studies that use dynamic analysis approaches to detect or fix performance bugs (especially for data loss issues). This result is align with expectation because dynamic analysis involves executing apps in a controlled environment and unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). Table 3 shows that three dynamic tools are designed to address data loss issues in the context of activity restarts associated with sophisticated lifecycle methods, which are challenging to be detected statically. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
Deep Learning. With the increasing popularity of deep learning (DL) techniques, we are surprised to find that only one paper adopted DL techniques for performance analysis. Specifically, Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) observed runtime features dynamically and leveraged DL models to locate energy anomalies. Although the number of DL-related studies is relatively small, DL techniques have been widely employed in various software engineering tasks, such as code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ). It is promising to see a garnered considerable research attention to employ ML/DL techniques to thwart Android performance issues.
Answers to RQ1.2: The majority of existing approaches primarily utilize static analysis (70.00%), with dynamic analysis (40.00%) coming next, and deep learning (10.00%). Only 10 out of the examined papers (20.83%) offer publicly accessible automated tools, highlighting a significant accessibility gap. Only 7 factors (33.33%) are addressed by existing automated tools, indicating incomplete coverage and potential oversight. Most tools prioritize issue detection over resolution, indicating a need for proactive problem-solving.
Table 4. A list of datasets used for Android mobile performance analysis.
Reference Year # APP Type Target Factor Link
Guo et al. (Guo2022DetectingAF, ) 2022 66 Real-World Activity  (Guo2022DetectingAFTool, )
Chen et al. (Chen2023TransparentRC, ) 2023 27 Real-World Activity  (Chen2023TransparentRCTool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Activity Resource  (Lu2022DetectingRUTool, )
Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) 2020 9 Micro Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro AsyncTask  (Pan2020StaticACTool, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Database Operations Loop  (Lyu2018RemoveRFTool, )
Song et al. (Song2019ServDroidDS, ) 2019 45 Real-World Service  (Song2019ServDroidDSTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World UI Rendering View Inefficient Method Hashmap  (Habchi2019TheROData, )
2.2.3. Result - Dataset
To evaluate the aforementioned automatic tools, researchers have proposed several datasets targeting unique performance factors. After examining all the public links or websites within the 48 identified studies, we were surprised to find only 9 (18.75%) unique open-sourced datasets, as shown in Table 4. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. Additionally, the number of factors covered by existing datasets is quite limited, w.r.t 11 factors in total. The shortage of dataset suggests an urgent need of comprehensive dataset construction. Moreover, there are a limited number (i.e., 89 apps per dataset on average) of apps in each dataset, making it unrealistic to cover all potential performance issues. In other words, even though some tools might effectively detect issues, they would still suffer from soundness issues due to the existence of false negatives. Overall, more effort should be dedicated to creating more representative and comprehensive datasets.
Answers to RQ1.3: Available datasets covers only 11 out of 21 performance factors, highlighting a significant gap in tool evaluation. Existing datasets lack the representation needed to cover all performance issues adequately.
3. A Large-scale Exploration of Real-world Apps
In RQ1 (Section 2), we systematically summarised the Android performance issues in academia. However, there may still be gaps between cutting-edge studies and real-world scenarios. Thus, we conducted an explanatory study by examining real-world performance issues discussions. Specifically, we collected reported performance issues from three popular platforms of Android ecosystem: Google Play, Stack Overflow, and GitHub. We 1) reviewed user feedback on Google Play to assess the impact on end-users; 2) explored the discussions on Stack Overflow to gain insights into common issues, discussions, and solutions shared within the developer community; and 3) analysed issues and commit histories on GitHub to understand the challenges developers are facing. While detailing research gaps, we also present code patterns that we believe could inspire future works, contributing to our broader community.
Figure 3. The process of real-world data collection.
3.1. Data Collection
Fig. 3 shows our three-step pipeline to obtain real-world performance issues, discussions, and solutions. Briefly, we first crawl user reviews from the Google Play Store, question threads from Stack Overflow, and issue-commit pairs from GitHub. Next, we filtered collected information by keywords obtained from RQ1. Lastly, we manually verified real-world performance issues, discussions, and their common patterns.
3.1.1. Data Crawling
To understand real-world performance issues, we collect data from three popular platforms in the Android development ecosystem as below.
[Google Play - User Reviews.] The Google Play Store is the biggest Android app store in the world. It constitutes an essential component of the Android ecosystem, facilitating the distribution and discovery of apps for millions of users worldwide. Users frequently complain about the performance issues of the Android apps by leaving reviews in the Google Play Store, providing useful information for researchers (Liu2014CharacterizingAD, ). We collected user reviews from the apps on Google Play based on the AppBrain popularity ranking [^2] AppBrain website https://www.appbrain.com/stats/google-play-rankings. To hold the representativeness, we collected the top 200 apps in five monetary categories (top free, top paid, top grossing, top new free, top new paid) across 30 countries, totaling 7,681 unique apps. We used the Google Play scraper library333google-play-scraper: https://pypi.org/project/google-play-scraper/ and successfully collected 909,430 user reviews from these apps. Please note that we excluded user reviews which only have a rating score (1-5) without comments. Intuitively, performance issues are embedded in the negative user reviews, so we used the sentiment analysis model SiEBERT (hartmann2023, ) to identify negative reviews. This model is a fine-tuned checkpoint of RoBERTa (Liu2019RoBERTaAR, ), with fine-tuning datasets from diverse text sources such as reviews and tweets. As a result, we roughly obtained 60,684 negative user reviews for further process.
[Stack Overflow - Question Threads.] Stack Overflow is a popular Q&A website for programmers and developers. It plays a crucial role in the Android ecosystem by providing a platform for knowledge sharing, problem-solving, and community support (Wu2022RetrievingAK, ; Huang2018APIMR, ). Thus, we explored the performance issues developers consider and discuss in the development process. Each question threads on Stack Overflow normally reflects one real-world programming challenges, and the follow-up discussion in the thread provide additional thoughts, insights, and potential solutions. To understand developers’ opinions on Android performance issues, we crawled all question threads labeled with “android” and containing accepted answers, resulting in 448,977 questions. We further narrowed down to 360,372 threads containing the code tag for further refinement.
[Github - Issues & Commits.] GitHub is a critical platform in the Android ecosystem. Issues and commits are essential tools on GitHub that help manage the development process in Android development projects. Issues provide a structured way to handle reported performance issues, while commits ensure a clear and traceable history of code changes as solutions. Therefore, issues and commits can provide valuable insights about how developers actually solve those issues in the maintenance process. We used F-Droid (FDroid, ), an open-source Android app repository, to access each app’s corresponding GitHub repository. Here, we chose F-Droid because it is way more easier to inspect code patterns in open-source apps compared to closed-source ones. To gather representative Android apps, we crawled 17 categories of open-source Android apps and their GitHub links from F-Droid. Based on these links, we successfully identify 4,228 Android apps and their GitHub repositories, containing 35,278 issues and 550,973 commits.
3.1.2. Keyword Filtering
As we focus on Android performance issues, we adapted a keyword-based filtering to select relevant data only. Alongside of the systematical literature review of 48 academic research papers in Section 2.1.4, we also snowballed a list of keywords about Android performance issues. The keyword list includes terms like “energy waste”, “response time”, “responsive”, and so on. As a previously unknown performance issue might be caused by an existing factor and vice versa, keywords from existing literature do not restrict us from discovering new performance issues and new factors. Using these keywords, we filtered 60,684 user reviews to 132, 360,372 Stack Overflow discussions to 773, 35,278 GitHub issues down to 297, and 550,973 GitHub commits to 645.
3.1.3. Manual Checking.
After filtering data by performance-related keywords, we notice some data entries are false positive, meaning they contained keywords but were not actually related to performance issues or their root causes. Therefore, we further conducted a manual verification to ensure the data quality and relevance.
User Reviews. We removed those unrelated to performance issues, such as criticisms of developers. For example, “The app works well, but only supports SMBv1, which has been deprecated and known to be very insecure. There have been numerous complaints about this but the author has been stubbornly unresponsive about it…”
Question Threads. For Stack Overflow, we excluded question threads that are wrongly labelled on the code tag and those low quality threads that their negative votes outnumber positive votes.
GitHub - Issues. We first checked and removed issues from inactive (dead) projects, such as those where developers do not respond or have explicitly stated that the project is no longer maintained. Additionally, we excluded issues that do not involve performance issues, such as “Show me … ports mac address host name mDNS response time, perhaps option to ping again and option to traceroute.
GitHub - Commits. We manually checked all remaining commits, and removed those without clear commit messages. Additionally, we excluded commits with multiple changes, such as fixing various bugs or features, where performance was only a minor part, making it difficult to identify the root causes.
Above all, we obtained 104 user reviews, 332 stack overflow discussions, 78 GitHub issues, and 289 GitHub commits. Two of authors performed the aforementioned manual checking individually, for any disagreement, they discussed and reached the same answer. For persisting disagreement, a senior author will then involve and decide. We utilized Cohen’s kappa (Landis1977AnAO, ) to assess inter-rater agreement. The results indicate a Cohen’s kappa of 0.953 for user reviews, 0.897 for commits, 0.923 for GitHub issues, and 0.885 for Stack Overflow posts, all of which signify substantial agreement. We then manually identified the Android performance issues in these data and present our findings in the next section.
[…]
Figure 4. Distribution of performance issues and factors in user reviews, stack overflow, GitHub issues, and GitHub commits. Other Factors: Polling, Syntax, Hasmap, Cookies, Override, Thread Pool, Cache, Copy Text, Background Thread, HTTP Request.
3.2. Result Analysis
Fig. 4 illustrates the distribution of performance issues and factors across the three data sources. Overall, we identified six categories of performance issues (including one previously unseen) and 35 factors (including 19 previously unseen) from the three real-world data sources. Next, we analyzed our findings from the perspectives of User Reviews, Stack Overflow, and GitHub Issues & Commits.
3.2.1. Performance Issues and Factors in Real-world
We first identified performance issues in real-world apps and highlight the gaps between these issues and those discussed in research communities.
[Google Play - User Reviews.] User reviews typically focus on how performance issues negatively impact their experience without delving into the specific causes. Some user reviews mention more than one performance issue, so the percentages add up to more than 100%. We totally identified six performance issues, five of which were known from the literature review, while one performance issue was previously unseen. Responsiveness issues, such as slow responsiveness or UI freezes, have the most direct impact on user experience, appearing 79 times (75.96%). For example, one comment says: “… the software appears to be sluggish in map updating and does not properly locate my vehicle… I am hopeful these issues will be resolved in the future but if not then I will stop discontinue my premium subscription service and find another solution.” This indicates that unresolved performance issues can lead to users canceling subscriptions, causing companies to lose market share. Energy Consumption is also critical, as inefficient app power consumption can quickly drain batteries. These issues appeared 25 times (24.04%). E.g., one representative comment states: “severe battery drain on s10, causing 1% loss every 3 minutes. drain immediately disappeared after uninstall, back to 1% loss every 8-9 minutes.” This shows that performance issues can lead to users uninstalling the app. Additionally, a performance issue not covered in existing literature is Data Consumption. E.g., one comment mentions: “.. Still cant even disable it, much less delete it…And worst of all, it still runs, using up data, memory and running down battery. Shame on You XXX…”
[Stack Overflow - Question Threads.] From the 332 question threads, we identified five categories of performance issues, all of which are already mentioned in the literature review (Section 2). Developers most frequently discussed Memory Consumption (80.72%), which is mainly caused by Image factor. However, the prevalence of Memory Consumption only ranks forth in existing research papers according to Fig. 2. Additionally, while Responsiveness is the most concerning issue for end users, it’s the second most frequent topic in developer discussions. This discrepancy indicates that the performance issues developers focus on, those users are concerned about, and the ones researchers prioritize, do not align. Moreover, we identified a total of 30 factors contributing to these performance issues, and only 16 factors appeared in the literature review. Among the 14 newly discovered factors, the most frequent one is File Operation, which typically involves inefficiently downloading, uploading, or saving large files, leading to memory shortages. The next most frequent previously unseen factor is Data Type, which often lead to Memory Consumption and Responsiveness issues.
[Github - Issues & Commits.] Among the 289 commits, we identified four categories of performance issues, all of which are already mentioned in the literature review (Section 2). The most frequent performance issue is Memory Consumption (84.43%), followed by Responsiveness (11.07%). This distribution closely resembles that found in Stack Overflow discussions, indicating that developers are most concerned with memory consumption. In addition, we identified 22 factors, 11 of which are previously unseen. For example, the factor Parameter can lead to data loss or memory consumption. In Commit  [^4] https://github.com/ente-io/ente/commit/0165b7569b4acee1699de59c7c6ec1a8b91a7b4f, the change was to remove the fileUrl from console.log parameters to prevent a memory leak.
Regarding GitHub issues, we identified four performance issues and nine factors (including three unseen factors). The most frequently discussed issue is Data Loss (51.28%), followed by Memory Consumption (38.46%) and Responsiveness (8.97%). Interestingly, Data Loss is the most common issue in GitHub Issues, while Memory Consumption is the most frequently addressed in commits. This discrepancy indicates that the issues reported in GitHub issues are not always the ones developers prioritize for fixing.
Answers to RQ2.1 We identified an unseen performance issue related to Data Consumption and 19 previously unrecognized factors such as Data Type, File Operation, and Object Creation, highlighting gaps in research coverage. The primary concerns of researchers, developers, and users are not aligned: researchers mainly focus on energy consumption, developers prioritize memory consumption, and users are most concerned with responsiveness.
3.2.2. Common Code Patterns in Real-World
After discovering numerous new factors, we then wanted to delve deeper into the prevalent code patterns associated with these issues. Through analysis of Stack Overflow discussions, GitHub issues, and commits, we aim to identify the performance issues commonly faced by developers during development and examine their code patterns. In total, we summarized six common patterns as follows:
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API order, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.
1const CopyTextOption = ({bottomSheetId, postMessage, sourceScreen}: Props) => {
2 const handleCopyText = useCallback(async () => {
3 ...
4 await dismissBottomSheet(bottomSheetId);
5 Clipboard.setString(postMessage);
6 };};
Listing 1: Code Example of API Order Misuse
Unreleased References Another significant pattern can be characterized by the failure to release references (or have them garbage collected). For example, if a Fragment is held as an implicit reference, it will not be garbage collected until the object holding the reference completes its execution. Initializing objects in complex Activity lifecycle methods can also result in unreleased references, ending up with memory leak. Listing 2 demonstrates a code snippet from Stack Overflow  [^5] https://stackoverflow.com/questions/56102382, where a singleton object holds an mContext reference (line 2). Here, mContext is a layer that stands behind its component (Activity, Service, etc.), and its lifecycle is much shorter than that of the singleton object. This prevents it from being garbage collected, leading to a memory leak.
1public class MySingleton() {
2 Private Context mContext;
3 Private static MySingleton mInstance;
4 ...
5 public void myMethod(Context context)
6 {this.mContext = context;...}}
Listing 2: Code Example of Unreleased Reference
Redundant Objects. The creation of redundant objects is a common pattern that causes performance issues. Developers often mistakenly create new instances repeatedly instead of referring to existing ones, leading to endless memory consumption. Additionally, when two classes are mutually dependent and create instances of each other during initialization, it results in endless object creation. Moreover, creating new objects with the same functionality can also unnecessarily degrades performance. Listing 3 demonstrates a code snippet from Stack Overflow  [^6] https://stackoverflow.com/questions/20942671, where every time setTypeface (line 2) is called, it creates a new Typeface object (line 4, 6, 8). This leads to increased memory usage, app slowdowns, and eventual out-of-memory crashes.
1public class MyTextView extends TextView {
2 public void setTypeface(Typeface tf, int style) {
3 if (style == Typeface.BOLD) {
4 super.setTypeface(Typeface.createFromAsset(...));
5 } else if (style == Typeface.ITALIC) {
6 super.setTypeface(Typeface.createFromAsset(...));
7 } else {
8 super.setTypeface(Typeface.createFromAsset(...));
9 }}}
Listing 3: Code Example of Redundant Objects
Large-Scale Data. Large-scale data is also a culprit for performance issues. Loading large images without proper caching can cause excessive memory usage and slowdowns. Uploading large files without separating them into smaller chunks can lead to high memory consumption. As shown in a code snippet from Stack Overflow  [^7] https://stackoverflow.com/questions/8820837, where the buffer size is the same as the uploaded file. If the uploaded file exceeds the maximum available memory, it will directly cause an out-of-memory error.
1 ...
2 int length = 0;
3 newFile.createNewFile();
4 InputStream inputStream = ctx.getAssets().open("myBigFile.apk");
5 FileOutputStream fOutputStream = new FileOutputStream(newFile);
6 byte[] buffer = new byte[inputStream.available()];
7 while ((length = inputStream.read(buffer)) > 0) {
8 fOutputStream.write(buffer, 0, length);
9 }
Listing 4: Code Example of Large-Scale Data
UI Operations. UI is another key factor that can affect performance, encompassing issues such as repeatedly drawing the same content and UI layout misconfiguration. For example, in the code snippet 5 from Stack Overflow  [^8] https://stackoverflow.com/questions/42944376, the ‘for’ loop in ‘onDraw’ (line 5) redraws the same circles every time when onDraw in invoked, even though the UI content doesn’t change. This causes the application to slow down and increases the computational load.
1public class ConcentricCircularView extends View {...
2 protected void onDraw(Canvas canvas) {
3 super.onDraw(canvas);
4 ...
5 for (int i=0;i<10;i++){...}}}
Listing 5: Drawing the Same Content in the Loop
Other Patterns. Besides the previously mentioned patterns, there are several other types of patterns to be aware of, such as scheduling a large number of Runnables to execute simultaneously, accessing CookieManager in the main thread, etc.
Answers to RQ2.2 We have meticulously identified six common code patterns that contribute to performance issues, which encompass API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns.
Figure 5. Performance issues and corresponding factors from literature review and real-world data. Colors represent the number of automated tools capable of analyzing each factor (0-5), with darker shades indicating more tools. White indicates no factor-to-performance relationship.
4. Future Trends
In this study, we summrized the relevant research on Android performance analysis and illuminated real-world challenges in Sections 2 and 3. Drawing upon these findings, we now present discussion points regarding the research and practice of Android performance analysis, aiming to guide future researchers in this field.
4.1. Targeted Objective Imbalance
Through our literature review (Section 2) and real-world exploration (Section 3), we identified a total of six types of performance issues and 40 factors, as shown in Fig. 5, where colors and white distinguish whether there is a relationship between factors and performance issues. We observed that, apart from the performance issue “Data Consumption”, all the other five performance issues involve corresponding factors, suggesting data consumption is an important but untouched topic.
In addition, of the 40 discovered factors, 21 have been analyzed in the literature, while the remaining 19 (almost half) have not been examined by our researchers, highlighting a significant research gap. Moreover, it’s intriguing to note that there are 5 factors (namely, Multi-Threads, Obsolete Task, Redundant Frames, SQL Statement, and Wakelock) that solely appear in the literature without a corresponding issue or complaint in real-world scenarios. This suggests that these factors may stem more from the assumptions of academic researchers rather than being genuine concerns of users or developers in practical settings.
The presence of this imbalance underscores the need for further research into issues genuinely concerning users or developers. Indeed, more effort should be put into addressing the untargeted real-world factors, aiming to bridge the existing gap between identified factors and practical concerns.
4.2. Automatic Performance Analysis Tools are Still in an Early Stage
Our review also revealed that only 7 out of 40 factors (17.50%) are covered by state-of-the-art (SOTA) tools. We utilize Fig. 5 to further highlight this significant research gap, where Activity is the most targeted factor, followed by Resource and AsyncTask, and finally Service, Database Operation, Image, and Loop. Also, the majority of these tools are focused on issue detection, while only 3 of them (30.00%) are dedicated to fixing such issues. Although static, dynamic, and DL techniques have been demonstrated to be effective at analyzing performance issues in a series of research experiments, applying these approaches to reveal performance issues holistically in practice remains unsolved. Future research should not only focus on proposing a comprehensive framework to tackle many realistic challenges, such as API misuse, redundant objects, etc, which may require knowledge of code patterns as presented in Section 3.2.2, but also on finding solutions for fixing problems.
In addition, the adopted techniques are also highly imbalanced (most existing tools use traditional static or dynamic analysis, with only one work (Jabbarvand2020AutomatedCO, ) leveraging DL methods). The lack of variety in technique selection hinders the soundness, as static approaches are not proficient in handling complicated code features (e.g., reflection (sun2021taming, ), obfuscation (samhi2022jucify, ), and hardening (Schloegel2022LokiHC, )), and it’s non-trivial for dynamic tools to achieve satisfactory results to trigger sufficient paths in practice. Thus, developing effective tools against performance issues will be a noteworthy research topic in the future. In fact, DL has made considerable achievements in software engineering (Wang2023MachineDeepLF, ; Feng2020CodeBERTAP, ; Gu2016DeepAL, ; Hu2018DeepCC, ), and these advanced techniques have demonstrated powerful capabilities in solving many complex tasks (i.e. code analysis (White2016DeepLC, ; Jebnoun2021ClonesID, ; Zhao2018DeepSimDL, ) and code generation (Lu2021CodeXGLUEAM, ; Wang2022TestDrivenML, )). Thus, it is promising to apply advanced DL techniques to assist in solving realistic and specific performance issues. For example, convert code exhibiting the Unreleased References pattern mentioned in Section 3.2.2 into program analysis graphs (e.g., CFG and DFG). Use these graphs as features to train or finetune a GNN-based model (Zhou2018GraphNN, ; Xu2018HowPA, ; Velickovic2017GraphAN, ) to identify this anti-pattern. Leverage the network’s interpretability (Ying2019GNNExplainerGE, ; Yuan2021OnEO, ) to pinpoint problematic code segments and suggest optimizations, such as adding code to release references.
4.3. Constructing Benchmark Dataset is an Urgent Need
Similar to the shortage of tools, current datasets are too sporadic to reflect real performance issues, evidenced by the fact that only 11 (27.50%) of the 40 factors are covered by existing datasets. Even among the 11 factors, the data distribution is highly imbalanced. The number of activity samples is much larger than that of the other factors, which are more severe from a real-world perspective. In addition, the number of available data in each dataset is very small (usually only a few dozen examples), which is insufficient to support a fair evaluation of automated tools. Moreover, another major issue is the limited reproducibility. Our review indicates that only a small number of primary studies share their datasets, making it difficult for future researchers to validate and evaluate tools. Furthermore, these datasets lack real-time updates for the latest app, thus containing outdated samples.
Thus, it is urgent to construct up-to-date, comprehensive benchmark datasets to support fair comparisons, covering as many real-world cases as possible. The code examples we summarized from the exploratory study can all be used to construct a comprehensive dataset, covering 35 factors. Another possible solution involves leveraging the findings of our exploratory study, wherein we identify six specific code patterns from real-world scenarios. These patterns can act as seeds for mutation tools, enabling the generation of additional data to construct datasets.
5. THREATS TO VALIDITY
Our study has several factors that could affect the validity of our findings in terms of literature review and large-scale exploration.
For literature review, despite using a well-established methodology (Kitchenham2004ProceduresFP, ), we cannot guarantee that our empirical results cover all relevant works. To locate relevant studies, we formulated search strings from both Android and performance issues, as described in Section 2.1.2, and used five well-known electronic databases. To minimize the risk of missing relevant works, we conducted backward and forward snowballing and cross-checked our results.
For real-world data collection, the data filtering process may miss some relevant information. This is because the sentiment classification model (hartmann2023, ) used to filter user reviews may not guarantee 100% accuracy in identifying negative reviews. Additionally, the Google Play scraper might fail to capture all user reviews. Lastly, there could be biases in our manual checking process. To minimize subjective differences, two authors independently checked the filtered data. In cases of disagreement, they discussed and resolved whether to include the data.
6. Conclusion
In this paper, we conducted a comprehensive empirical study to reveal the current state of Android performance analysis and the research gap between real-world challenges and existing techniques. Based on our in-depth analysis of 48 relevant studies and real-world issues across three major platforms (Google Play, Stack Overflow, GitHub), we discovered that (1) There are 6 types of performance issues and corresponding 40 factors; (2) Performance issues are sophisticated, with a single issue potentially stemming from multiple factors and a single factor causing multiple issues; (3) The root causes of data consumption issue are not clear; (4) The primary concerns of researchers, developers, and users are not aligned; (5) The majority of existing approaches primarily utilize static analysis; (6) There is a significant gap in existing tools’ ability to address all factors causing performance issues; (7) There is an urgent need to construct benchmark dataset to evaluate automatic tools; (8) DL-based and LLM-based approaches have shown promising results in addressing software engineering problems, suggesting their potential for tackling Android performance issues as well.